angelinagolban:

The temperature at a mountain base camp was −4 degrees Celsius on Saturday. Sunday morning, the temperature was 2 degrees Celsius lower than it was on Saturday. By Sunday evening, the temperature was 1 degree Celsius lower than it was that morning. What was the temperature at the base camp Sunday evening?

justus:

The equation you can use for this problem is -4 - 2 - 1 -4 - 2 - 1= -6 - 1=

justus:

Now all you have to do is -6-1=

angelinagolban:

ok thx

justus:

You're welcome.

angelinagolban:

ok im working it out

justus:

Good luck